News by Date
Open and close menu tab
News by Category
Open and close menu tab
Press Resources
Open and close menu tab
Web Shortcuts
Open and close menu tab
Operations CXO Status Report

Friday 28 May 2004 8.30am EDT

During the last week the observing schedule was replanned to include an observation of SGR 1806-20 that was approved as a fast 3-day turn-around Target Of Opportunity (TOO) on 25 May. Loads were replanned and uplinked on May 26 for observing May 27. Observations of RCrADarkCloud, LBQS 0010-0012 and CL0853+5759 were impacted by the TOO and will be rescheduled for a later load.

A real-time procedure was completed on 24 May as part of a diagnosis of a stuck-on heater located on the Optical Bench Assembly (OBA) in Zone 50. The procedure turned off power to a set of survival heaters, including the one in Zone 50. After seeing no thermal response, the power was turned back on. The results indicated that the survival heater is functioning as expected (i.e. it is not stuck on). Additional work will proceed to trace the fault to the primary or redundant operational heater.

The schedule of targets for the next week is shown below and includes 2 observations of the Moon.

------------------------------------------
A2069                  ACIS-I       May 30
Radiation Belts                     May 31
36 Oph                 HRC-S/LETG   Jun 1
Moon (2 obs)           HRC-I        Jun 2
Circinus Galaxy        ACIS-S/HETG
2QZ J003058.2-2756     ACIS-S       Jun 3
Radiation Belts
MACS0140.0-0555        ACIS-I
0454-463               ACIS-S
4U 1624-49             ACIS-S/HETG  Jun 4
HD101584               ACIS-S
NGC 5347               ACIS-S       Jun 5
HR 5875                ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports